The probability that a man will live 5 years more is 3/4 and the probability that his wife will live 5 years more is 1/3. What is the probablity that both will stil live 5years more?

Thank you for your help..=)

Answer by Mathtut(3670)

(3/4)*(1/3)=3/12=1/4 or .25(100)=25% chance that both will live 5 more years.